A stream of liquid cyclopentane flows at a rate of 0.35lbm/s

SpecificA stream of liquid cyclopentane lows at a rate of 0.35 1bm/s into an Gravity and initially empty container with a volume of 2 ft3, placed on a scale Densitycalbrated to read weight in lbe The process takes place at sea level. When only the empty container was on the scale, the scale reading was 2.50 lb. Estimate the reading 10 seconds after the flow starts. 2.85 6.00 10.00 If the container were being filed at the top of a mountain, would the mass of liquid in the container after 10 seconds be exactly the same as it was at sea level?

Yes No Would the weight of the liquid in the container after 10 seconds be exactly the same on top of the mountain as it was at sea level? Yes No Look up the specific gravity of liquid cyclopentane, and calculate the density of this fluid in Ib/t 1.00 0.745 46.5 62.4

On a scratch sheet, write a dimensional equation for the time (seconds) it will take to fill the container. Then check the corect numerical expression for this quantity 2x0.35/46.5 2x46.5/0.35 0.35/ 2x46.5) D 46.5/(2x3)
Answer #1

given, mass flowrate,m=0.35lbm/s; Vc =2 ft3; initial weight(reading), Wo = 2.5 lbf; gravitational acceleration, g = 32.2 ft/s2

mass of solution accumulated in container after 10 seconds, M= t*v = 10*0.35 = 3.5 lbm

new reading, W = Wo + M = 2.5 + 3.5 = 6 lbf

mass of any object is independent of location; answer: YES

Since, the gravitational force will be slightly weaker on the top of mountain hence the reading will reduce; answer:NO

specific gravity of cyclopentane, s= 0.7454

density of cyclopentane,d =0.7454*62.428 =46.5 lb/f3

volumetric flow rate,v = m/d = 0.35/46.5

time taken to completely, T = Vc/v